Suppose the economy is operating at equilibrium, with Y0 5 1,000. If the government undertakes a fiscal change whereby the tax rate, t , increases by .05 and government spending increases by 50, will the budget surplus go up or down? Why?
1
Expert's answer
2021-08-16T14:47:07-0400

Budget surplus "(BS)=tY-G-TR"

Tax increases by 0.05 as government spending increases by 50

Therefore

"t_2-t_1=0.05\\and\\G_2-G_2=50\\so\\∆BS=(t_2-t_1)Y-(G_2-G_1)\\=(0.05×51000)-50\\=2550-50\\=2500"

Change in budget surplus is 2500

The budget surplus will go up.

Increase in taxes and government spending raises the revenue.